15. Convert (256 — i512)ohms to polar coordinates.
Answer #1

The polar form is written as

A\angle \theta

where

A = sqrt (2562 + 5122)

A = 572.43

\theta = arctan (-512/256)

\theta = -63.43 so

polar form is

572.43 \angle-63.43